# ReservoirTTA: Prolonged Test-time Adaptation for Evolving and Recurring Domains

### v1

@@ Appendix F Model Reservoir Initialization
In this section, we explore strategies for initializing a new reservoir model when a new domain is detected.
As shown in Figure 12, we evaluate two initialization approaches: (1) initializing with source model parameters and (2) initializing with reservoir parameters that maximize mutual information (Equation 11) on the current batch. We evaluate these on CIFAR100-C under recurring CSC and CDC settings (20 recurrences) using ResNeXt-29, and on CCC with ResNet-50 and ImageNet-C under recurring CSC and CDC settings (20 recurrences) using a ViT-B-16. Our results show that the optimal strategy depends on both the method and the backbone. For convolutional networks, TENT and ETA benefit from source initialization—likely because, without a regularization module, drifting too far from the source degrades performance—while EATA is largely unaffected. Conversely, for ViT-B-16, initializing via Equation 11 significantly reduces error rates after 20 recurrences for TENT, ETA, and EATA, suggesting that the transformer is less prone to parameter drift and can better leverage the most similar model in the reservoir, whereas source initialization may overwrite valuable acquired knowledge.
@@ Appendix G Baselines
Table 8: Hyperparameter settings for TTA experiments. Following RobustBench, we use WideResNet-28 [42] for CIFAR-10-C, ResNeXt-29 [40] for CIFAR-100-C, and ResNet-50 [9] for ImageNet-C. We also evaluate on ViT-B-16 [5] and ResNet-50 with GroupNorm [39] for ImageNet-C. ReservoirTTA-specific settings are provided below.
@@
  • PeTTA [11]: Explicitly controls parameter drift during test-time adaptation by dynamically adjusting the update and regularization parameters, thereby preventing model collapse and sustaining performance over extended, recurring testing scenarios.
@@ Appendix H Implementation Details
Following the RobustBench [^3] https://github.com/RobustBench/robustbench [10] (MIT license), we employ WideResNet-28 [42] on CIFAR-10-C, ResNeXt-29 [40] on CIFAR-100-C, and ResNet-50 [9] on ImageNet-C. We also evaluate on ViT-B-16 [5] and ResNet-50 with GroupNorm [39] for ImageNet-C to further assess generalizability.
For CIFAR-10-C and CIFAR-100-C, TTA baselines (except SAR [25]) are optimized with the Adam optimizer [15] using a learning rate of $1\times 10^{-3}$, a universal $\beta=(0.9,0.999)$, and a batch size of 200, whereas SAR employs SGD [28]. For ImageNet-C, models are adapted with SGD at a batch size of 64 and a learning rate of $2.5\times 10^{-4}$ (adjusted to $1\times 10^{-4}$ for ViT-B-16 in the CCC setting). For ReservoirTTA, we configure the system with a maximum of $K^{\text{max}}=16$ reservoirs, determine the threshold $\tau$ using 2000 source examples, and update centroids with AdamW [20] at a learning rate of $1\times 10^{-4}$. Table 8 summarizes these settings. Note CIFAR10-C, CIFAR100-C, and ImageNet-C are publicly available online [^4] https://github.com/hendrycks/robustness (Apache-2.0 license). CCC is also provided by Rdumb paper [^5] https://github.com/oripress/CCC [27] (MIT license). All of our experiments have been conducted on a single NVIDIA A100 Tensor Core GPU (80 GB VRAM) using an internal cluster.
@@ Appendix I Additional Ablation Studies
Figure 15: Sensitivity study to the batch size $B$ on CIFAR-100-C under recurring CSC and CDC settings.
@@
Impact of ReservoirTTA on Regularization Weight Sensitivity of EATA and ROID. We analyze how ReservoirTTA affects the optimal hyperparameter $\alpha$ of ROID, which controls the momentum in weight ensembling with the source model. We also examine its impact on Fisher weight $\lambda$, which regulates the distance to the source model parameters in EATA. As shown in Figure 13, ReservoirTTA reduces the sensitivity of both EATA and ROID to these hyperparameters on CIFAR100-C recurring CSC and CDC settings. Furthermore, we observe that the best results with ReservoirTTA are achieved at a higher $\alpha$ for ROID, shifting from 0.99 in ROID to 0.995 in ROID+ReservoirTTA. A similar trend is observed for EATA, where the optimal Fisher weight $\lambda$ decreases from 2000 in EATA to 1000 in EATA+ReservoirTTA. This effect can be attributed to the Model Reservoir, which exposes each model to a more stable distribution of domains over time, thereby reducing the need for strong regularization.
This effect is even more pronounced in ViTs. As shown in Figure 14, EATA achieves its best performance at $K^{\text{max}}$=1 with a regularization weight of 2000.0, whereas ReservoirTTA ($K^{\text{max}}$=16) reaches optimal performance with a regularization weight of 0. Interestingly, as $K^{\text{max}}$ increases, the need for regularization diminishes. For ROID, the trend differs. The momentum parameter $\alpha$ should be set to 1.0 for ViT to achieve the best results, suggesting that regularization is unnecessary for this architecture. Moreover, further increasing $K^{\text{max}}$ reduces the error rate by 0.6% on CCC, demonstrating the effectiveness of ReservoirTTA in this setting.
@@
Effect of Batch Size. As illustrated in Figure 15, we explore batch sizes from 8 to 200 for both EATA + ReservoirTTA and ROID* + ReservoirTTA on CIFAR-100-C under recurring CSC and CDC settings. We observe a consistent decrease in error rate with larger batches, which likely stems from more stable gradient estimates, reduced variance in style features, and more accurate domain assignments. Conversely, very small batches can lead to noisy updates and suboptimal domain clustering. Although a large batch is important for reliable performance, memory constraints and real-time needs may limit its size in practice.
Sensitivity to Various Sequence Orders. We examine how DPCore, CoLA, and ReservoirTTA respond to different test batch orders during domain discovery. As shown in Figure 16, we measure the number of detected domains after the first visit and across 20 visits in recurring CSC on CIFAR100-C using five seeds—each corresponding to a distinct batch order while keeping the domain sequence fixed. Our findings reveal that DPCore is highly sensitive to batch order, undermining its reliability. Although CoLA yields consistent detection across seeds, its number of detected domains continuously increases, reaching 160 models after 20 visits. In contrast, ReservoirTTA shows minimal sensitivity, with detected domains stabilizing near the ground truth of 15 after 20 visits. These results underscore the superiority of our style quantifier, online clustering, and model discovery mechanisms over those used in CoLA and DPCore.
@@
Figure 17 shows pseudo segmentation label visualization on the ACDC dataset [29]. The first two columns display outputs from TENT and BECoTTA, respectively, while the third column presents refined labels from BECoTTA combined with ReservoirTTA. The inclusion of ReservoirTTA yields more detailed and accurate segmentation labels, improving overall prediction quality.
@@
Can ReservoirTTA Help Tackle Temporally Correlated Test Streams? ReservoirTTA raises the question of whether it can tackle the challenge of a temporally correlated testing stream. As observed in Table 21—on CIFAR-10 $\rightarrow$ CIFAR-10-C under recurring continual CSC, with temporal correlation of image categories (Dirichlet coefficient = 0.1)— where we integrate ReservoirTTA with state-of-the-art TTA method, RoTTA [41], which is designed to address temporal correlations—yields similar gains. RoTTA tackles the instability caused by temporally correlated test samples by introducing mechanisms that stabilize the online adaptation process via smoothing the updates over time. In particular, RoTTA uses an exponential moving average (EMA)–based update to accumulate robust, long-term statistics that mitigate the bias introduced by correlated mini-batches. This principle is further enforced by using robust batch normalization (RBN) and by maintaining a memory bank through category-balanced sampling, which ensures a representative snapshot of the test distribution.
@@
Table 21: Average classification error rates (%) for WideResNet-28 on CIFAR-10 $\rightarrow$ CIFAR-10-C at severity level 5 under recurring continual CSC, with temporal correlation of image categories (Dirichlet coefficient = 0.1).
Recurring TTA visit $\xrightarrow{\hskip 227.62204pt}$
Method 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 Avg
Source 43.52 43.52
RoTTA 26.62 25.17 25.05 25.56 25.20 25.26 25.88 27.07 28.50 28.78 30.28 30.19 32.06 32.96 34.30 35.66 35.38 36.70 37.93 38.87 30.37
+ ReservoirTTA 26.19 25.32 25.66 25.69 25.72 25.88 25.26 25.63 26.44 25.85 26.67 25.78 25.94 25.56 26.35 26.95 25.87 25.54 25.86 25.67 25.89 -4.48

### v3

@@ Appendix F Model Reservoir Initialization
In this section, we explore strategies for initializing a new model when a new domain is detected.
MI vs. Source Weight Initialization. As shown in Figure˜13, we evaluate two initialization approaches: (1) initializing with source model parameters and (2) initializing with reservoir parameters that maximize mutual information (Equation˜11) on the current batch. We evaluate these on CIFAR100-C under recurring CSC and CDC settings (20 recurrences) using ResNeXt-29, and on CCC with ResNet-50 and ImageNet-C under recurring CSC and CDC settings (20 recurrences) using a ViT-B-16. Our results show that the optimal strategy depends on both the method and the backbone. For convolutional networks, TENT and ETA benefit from source initialization—likely because, without a regularization module, drifting too far from the source degrades performance—while EATA is largely unaffected. Conversely, for ViT-B-16, initializing via Equation˜11 significantly reduces error rates after 20 recurrences for TENT, ETA, and EATA, suggesting that the transformer is less prone to parameter drift and can better leverage the most similar model in the reservoir, whereas source initialization may overwrite valuable acquired knowledge.
MI vs Top-k Model Weight Ensembling Initialization. We assess the effect of top-$k$ ($k=3$) ensemble initialization strategy for EATA+ReservoirTTA. On CIFAR-100-C (CSC) with a CNN backbone, we observe that EATA remains unaffected by the choice of initialization, yielding nearly identical performance across source, top-$k$, and MI initialization, consistent with our previous analysis. In contrast, on ImageNet-C with a ViT backbone, top-$k$ initialization improves performance by 3.37% compared to source initialization, but degrades performance over MI initialization (–0.27%).
@@ Appendix G Baselines
  • PeTTA [12]: Explicitly controls parameter drift during test‐time adaptation by dynamically adjusting the update and regularization parameters, thereby preventing model collapse and sustaining performance over extended, recurring testing scenarios.
Table 10: Hyperparameter settings for TTA experiments. Following RobustBench, we use WideResNet-28 [45] for CIFAR-10-C, ResNeXt-29 [43] for CIFAR-100-C, and ResNet-50 [10] for ImageNet-C, DomainNet-126, and PACS. We also evaluate on ViT-B-16 [6] and ResNet-50 with GroupNorm [42] for ImageNet-C. ReservoirTTA-specific settings are provided below.
@@
DomainNet-126 (ResNet-50) SGD $2.5\times 10^{4}$ 128
PACS (ResNet-50) Adam $1\times 10^{-3}$ 64
@@ Appendix H Implementation Details
Following the RobustBench [^3] https://github.com/RobustBench/robustbench [11] (MIT license), we employ WideResNet-28 [45] on CIFAR-10-C, ResNeXt-29 [43] on CIFAR-100-C, and ResNet-50 [10] on ImageNet-C. We also evaluate on ViT-B-16 [6] and ResNet-50 with GroupNorm [42] for ImageNet-C to further assess generalizability. For CIFAR-10-C and CIFAR-100-C, TTA baselines (except SAR [27]) are optimized with the Adam optimizer [16] using a learning rate of $1\times 10^{-3}$, a universal $\beta=(0.9,0.999)$, and a batch size of 200, whereas SAR employs SGD [31]. For ImageNet-C, models are adapted with SGD at a batch size of 64 and a learning rate of $2.5\times 10^{-4}$ (adjusted to $1\times 10^{-4}$ for ViT-B-16 in the CCC setting). For ReservoirTTA, we configure the system with a maximum of $K^{\text{max}}=16$ reservoirs, determine the threshold $\tau$ using 2000 source examples, and update centroids with AdamW [22] at a learning rate of $1\times 10^{-4}$. Table˜10 summarizes these settings. Concerning DomainNet-126, and PACS, we optimize the models with SGD at $2.5\times 10^{-4}$ learning rate, and Adam at $1\times 10^{-3}$ learning rate, respectively. Note CIFAR10-C, CIFAR100-C, and ImageNet-C are publicly available online [^4] https://github.com/hendrycks/robustness (Apache-2.0 license). CCC is also provided by Rdumb paper [^5] https://github.com/oripress/CCC [30] (MIT license). Both DomainNet-126666https://ai.bu.edu/M3SDA/ and PACS [^7] https://huggingface.co/datasets/flwrlabs/pacs are publicly available. All experiments were run on a single NVIDIA A100 Tensor Core GPU (80 GB VRAM) on our internal cluster.
@@ Appendix I Additional Ablation Studies
Figure 16: Sensitivity study to the batch size $B$ on CIFAR-100-C under CSC and CDC settings.
@@
Impact of ReservoirTTA on Regularization Weight Sensitivity of EATA and ROID. We analyze how ReservoirTTA affects the optimal hyperparameter $\alpha$ of ROID, which controls the momentum in weight ensembling with the source model. We also examine its impact on Fisher weight $\lambda$, which regulates the distance to the source model parameters in EATA. As shown in Figure˜14, ReservoirTTA reduces the sensitivity of both EATA and ROID to these hyperparameters on CIFAR100-C recurring CSC and CDC settings. Furthermore, we observe that the best results with ReservoirTTA are achieved at a higher $\alpha$ for ROID, shifting from 0.99 in ROID to 0.995 in ROID+ReservoirTTA. A similar trend is observed for EATA, where the optimal Fisher weight $\lambda$ decreases from 2000 in EATA to 1000 in EATA+ReservoirTTA. This effect can be attributed to the Model Reservoir, which exposes each model to a more stable distribution of domains over time, thereby reducing the need for strong regularization.
This effect is even more pronounced in ViTs. As shown in Figure˜15, EATA achieves its best performance at $K^{\text{max}}$=1 with a regularization weight of 2000.0, whereas ReservoirTTA ($K^{\text{max}}$=16) reaches optimal performance with a regularization weight of 0. Interestingly, as $K^{\text{max}}$ increases, the need for regularization diminishes. For ROID, the trend differs. The momentum parameter $\alpha$ should be set to 1.0 for ViT to achieve the best results, suggesting that regularization is unnecessary for this architecture. Moreover, further increasing $K^{\text{max}}$ reduces the error rate by 0.6% on CCC, demonstrating the effectiveness of ReservoirTTA in this setting.
Table 11: Comparison of reset-based baselines and ReservoirTTA on CIFAR100-C (CSC). Error (%) at visits 1–20 and mean.
Visits 1–10
Method 1 2 3 4 5 6 7 8 9 10
RDumb 31.95 31.49 32.34 32.85 31.96 31.51 32.24 32.99 32.07 31.48
RDumb w/ VGG 31.11 31.13 31.17 31.20 31.19 31.16 31.20 31.21 31.20 31.21
EATA+ReservoirTTA 30.56 29.07 28.75 28.58 28.52 28.46 28.41 28.41 28.39 28.39
Visits 11–20
Method 11 12 13 14 15 16 17 18 19 20 Avg
RDumb 32.29 32.82 32.07 31.48 32.41 32.88 32.15 31.43 32.22 32.86 32.17
RDumb w/ VGG 31.12 31.17 31.16 31.17 31.11 31.17 31.17 31.16 31.20 31.19 31.17
EATA+ReservoirTTA 28.38 28.40 28.37 28.37 28.37 28.40 28.36 28.43 28.37 28.38 28.57
@@
Effect of Batch Size. As illustrated in Figure˜16, we explore batch sizes from 8 to 200 for both EATA + ReservoirTTA and ROID* + ReservoirTTA on CIFAR-100-C under recurring CSC and CDC settings. We observe a consistent decrease in error rate with larger batches, which likely stems from more stable gradient estimates, reduced variance in style features, and more accurate domain assignments. Conversely, very small batches can lead to noisy updates and suboptimal domain clustering. Although a large batch is important for reliable performance, memory constraints and real-time needs may limit its size in practice.
Sensitivity to Various Sequence Orders. We examine how DPCore, CoLA, and ReservoirTTA respond to different test batch orders during domain discovery. As shown in Figure˜17, we measure the number of detected domains after the first visit and across 20 visits in recurring CSC on CIFAR100-C using five seeds—each corresponding to a distinct batch order while keeping the domain sequence fixed. Our findings reveal that DPCore is highly sensitive to batch order, undermining its reliability. Although CoLA yields consistent detection across seeds, its number of detected domains continuously increases, reaching 160 models after 20 visits. In contrast, ReservoirTTA shows minimal sensitivity, with detected domains stabilizing near the ground truth of 15 after 20 visits. These results underscore the superiority of our style quantifier, online clustering, and model discovery mechanisms over those used in CoLA and DPCore.
Resetting to Source vs. Reusing Specialists. We analyze reset-based strategies to isolate the benefit of reusing adapted specialists versus simply resetting to source parameters. On CIFAR-100-C (CSC), blind periodic resets every 1,000 iterations (RDumb) yield an average error of 32.17%. Domain-aware resets that trigger only when the VGG-based detector flags a domain change (RDumb w/ VGG) improve to 31.17% ($1.0\%$). By contrast, ReservoirTTA+EATA reactivates the specialist previously adapted to a recurring domain—avoiding costly re-adaptation and forgetting—and achieves 28.57% ($2.60\%$ vs. domain-aware, $3.60\%$ vs. blind). See Table˜11 for the comparison results.
@@
Figure˜18 shows pseudo segmentation label visualization on the ACDC dataset [33]. The first two columns display outputs from TENT and BECoTTA, respectively, while the third column presents refined labels from BECoTTA combined with ReservoirTTA. The inclusion of ReservoirTTA yields more detailed and accurate segmentation labels, improving overall prediction quality.
@@
Additional Evaluation on Object-Level Style Shifts. To further assess the versatility of our framework beyond scene-level corruptions, we evaluate on datasets with pronounced object-level domain gaps, namely DomainNet-126 (126 classes) and PACS (7 classes). On DomainNet-126, we follow a CSC protocol by training a ResNet-50 on each source domain (real, painting, clipart, sketch), then adapt over a 20-cycle recurring sequence of the remaining three target domains. A similar protocol is applied to PACS, using photo, art painting, cartoon, and sketch as sources. As reported in Tables˜24 and 25, integrating ReservoirTTA with EATA consistently reduces error across both datasets, yielding average improvements of 0.79% on DomainNet-126 and 0.70% on PACS, with the largest gains observed when adapting from Real on DomainNet-126 (–1.60%) and from Cartoon on PACS (–1.93%). These results show that a framework motivated by scene-level shifts transfers robustly to object-level distribution changes.
Evaluation Under Gradual Shift. To assess the ability of our framework to handle gradual changes in corruption severity, we evaluate on a gradual domain shift stream constructed from CIFAR-100-C (CSC) using a ResNeXt-29 backbone. Each corruption type cycles through severity levels $1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1$ over 20 recurrences. Even under these subtle transitions, ReservoirTTA consistently improves over EATA. As shown in Table˜26, EATA averages $25.41\%$ error, while ReservoirTTA achieves $25.06\%$ with $K_{\max}=16$ ($0.35\%$) and $24.95\%$ with $K_{\max}=64$ (additional $0.10\%$). Thus, our method adapts to gradually evolving severities, not only abrupt domain changes.
Can ReservoirTTA Help Tackle Temporally Correlated Test Streams? ReservoirTTA raises the question of whether it can tackle the challenge of a temporally correlated testing stream. As observed in Table˜27—on CIFAR-10 $\rightarrow$ CIFAR-10-C under recurring continual CSC, with temporal correlation of image categories (Dirichlet coefficient = 0.1)— where we integrate ReservoirTTA with state-of-the-art TTA method, RoTTA [44], which is designed to address temporal correlations—yields similar gains. RoTTA tackles the instability caused by temporally correlated test samples by introducing mechanisms that stabilize the online adaptation process via smoothing the updates over time. In particular, RoTTA uses an exponential moving average (EMA)–based update to accumulate robust, long-term statistics that mitigate the bias introduced by correlated mini-batches. This principle is further enforced by using robust batch normalization (RBN) and by maintaining a memory bank through category-balanced sampling, which ensures a representative snapshot of the test distribution.
@@
Table 24: Comparison on DomainNet-126 under CSC setting (error rates %). Each source domain is trained separately; adaptation occurs over a recurring cycle of the remaining three domains repeated 20 times.
Source Domain real painting clipart sketch
Target Domain
clipart $\rightarrow$ painting
$\rightarrow$ sketch $(\times 20)$
real $\rightarrow$ sketch
$\rightarrow$ clipart $(\times 20)$
sketch $\rightarrow$ real
$\rightarrow$ painting $(\times 20)$
painting $\rightarrow$ real
$\rightarrow$ clipart $(\times 20)$
Avg
Source 45.16 41.57 49.52 45.33 45.40
EATA 38.49 33.49 38.34 33.34 35.92
+ReservoirTTA 36.89 32.25 38.65 32.74 35.13
$\Delta$ -1.60 -1.24 +0.31 0.60 -0.79
Table 25: Comparison on PACS under CSC setting (error rates %). Each source domain is trained separately; adaptation occurs over a recurring cycle of the remaining three domains repeated 20 times.
Source Domain photo art painting cartoon sketch
Target Domain
cartoon $\rightarrow$ art painting
$\rightarrow$ sketch $(\times 20)$
sketch $\rightarrow$ photo
$\rightarrow$ cartoon $(\times 20)$
photo $\rightarrow$ sketch
$\rightarrow$ art painting $(\times 20)$
cartoon $\rightarrow$ art painting
$\rightarrow$ photo $(\times 20)$
Avg
Source 56.97 32.21 23.71 72.06 46.24
EATA 38.22 22.80 18.82 37.68 29.38
+ReservoirTTA 37.79 22.55 18.63 35.75 28.68
$\Delta$ -0.43 -0.25 -0.19 -1.93 -0.70
Table 26: Performance under gradual shift on CIFAR-100-C (CSC, ResNeXt-29). Each corruption cycles severities $1\!\to\!2\!\to\!3\!\to\!4\!\to\!5\!\to\!4\!\to\!3\!\to\!2\!\to\!1$. Error (%) at visits 1–20 and mean.
Visits 1–10
Method 1 2 3 4 5 6 7 8 9 10
Source 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58
EATA 25.35 25.44 25.41 25.41 25.40 25.43 25.39 25.44 25.39 25.38
+ReservoirTTA ($K_{\max}=16$) 25.17 25.06 24.99 25.03 25.00 25.02 25.01 25.09 25.09 25.05
+ReservoirTTA ($K_{\max}=64$) 25.17 24.99 24.92 24.93 24.94 24.93 24.92 24.95 24.95 24.91
Visits 11–20
Method 11 12 13 14 15 16 17 18 19 20 Avg
Source 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58 33.58
EATA 25.45 25.42 25.44 25.44 25.43 25.42 25.42 25.43 25.38 25.40 25.41
+ReservoirTTA ($K_{\max}=16$) 25.07 25.08 25.05 25.10 25.07 25.08 25.09 25.05 25.06 25.07 25.06
+ReservoirTTA ($K_{\max}=64$) 24.94 24.91 24.94 24.94 24.94 24.93 24.97 24.97 24.97 24.96 24.95
Table 27: Average classification error(%) for WideResNet-28 on CIFAR-10 $\to$ CIFAR-10-C (severity 5) under recurring CSC with temporal correlation (Dirichlet = 0.1).
Visits 1–10
Method 1 2 3 4 5 6 7 8 9 10
Source 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52
RoTTA 26.62 25.17 25.05 25.56 25.20 25.26 25.88 27.07 28.50 28.78
+ReservoirTTA 26.19 25.32 25.66 25.69 25.72 25.88 25.26 25.63 26.44 25.85
Visits 11–20
Method 11 12 13 14 15 16 17 18 19 20 Avg
Source 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52 43.52
RoTTA 30.28 30.19 32.06 32.96 34.30 35.66 35.38 36.70 37.93 38.87 30.37
+ReservoirTTA 26.67 25.78 25.94 25.56 26.35 26.95 25.87 25.54 25.86 25.67 25.89 -4.48